# Nucleus and Postperihelion Activity of Interstellar Object 3I/ATLAS Observed by Hubble Space Telescope

### v1

@@ IV.2 Statistics of Similar Interstellar Interlopers
We attempted to evaluate the statistics of interstellar interlopers similar in size to 3I inside our solar system. Based on 1I/‘Oumuamua alone, D. Jewitt et al. (2017) and A. Do et al. (2018) derived a number density for such objects of $\sim\!0.1$-0.2 au-3 in the solar system. The subsequent discovery of the larger object 2I/Borisov was not a fluke but was in line with the extrapolated statistics (D. Jewitt et al., 2020). Recently, D. Z. Seligman et al. (2025) reported a number density of $\sim\!3\times 10^{-4}$ au-3 for 3I-like objects with $H_{V}=12.5$. However, since this absolute magnitude overestimated the nucleus size due to contamination from the coma, the real number density was underestimated. From Figure 8 in D. Z. Seligman et al. (2025), we estimated an updated number density of $\sim\!3\times 10^{-3}$ au-3 for 3I-sized interstellar objects as detected by the ATLAS survey. Given the discovery of 3I at a heliocentric distance of $r_{\rm H}\approx 4.5$ au, this implies a total number of $\sim\!1$ similar interstellar object within a sphere of this radius at the Sun at any instant on average. Therefore, the discovery of 3I is not unexpected.
@@
If extrapolated from the statistics solely based on 1I/‘Oumuamua using a differential size distribution index of -3, the total number of 3I-sized or larger interstellar objects inside the same volume of space around the Sun will be $\sim\!0.2$-0.3. The likelihood of detecting such an object is given by Poisson statistics to be $\sim\!0.2$-0.5, which drops to $\sim\!\left(1{\rm-}2\right)\%$ for a steeper size distribution of -4. However, these likelihood estimates should only be treated as lower limits because cometary activity undoubtedly enhances the detectability of 3I-like objects. We conclude that the discovery of 3I remains in qualitative agreement with the previous statistics of the interstellar object population. The arrival rate of 3I-like interstellar objects, estimated from the ratio of their total number within the 4.5 au radius sphere to the corresponding crossing timescale, $\sim\!2\times 10^{7}$ s, is $\sim\!1$ per year.
@@
Although all-sky surveys such as Pan-STARRS and Catalina have been operational for over a decade, only two other interstellar objects were previously discovered, both of which were visually fainter than 3I. Around its peak brightness, 3I was sufficiently bright that even visual comet hunters could have detected it had it passed through the inner solar system prior to the era of automated sky surveys. Here we simplistically estimated the likelihood that no 3I-like interstellar objects passed by between mid-1990s, the advent of modern sky surveys, and the discovery of 1I/‘Oumuamua in 2017. Using Poisson statistics for this $\sim\!20$-year interval, we found the probability of no detectable 3I-like objects to be $\sim\!10^{-13}$, which is unrealistically minuscule. While earlier surveys were undeniably less sensitive if compared to contemporaneous ones, it seems unlikely that they were less efficient by many orders of magnitude. Therefore, it appears highly probable that several 3I-like interstellar objects passed through the inner solar system undetected between the advent of automated sky surveys and the discovery of 1I/‘Oumuamua.
@@ IV.3 Dust Activity & Scattering Phase Function
Preperihelion continuum photometry from $r_{\rm H}=4.6$ au to 1.8 au revealed a steady brightening of 3I well described by an activity index $n=3.8\pm 0.3$ (D. Jewitt & J. Luu, 2025). Our results show that over similar ranges of heliocentric distance, 3I faded more rapidly on the outbound leg than it brightened on the inbound leg. [^4] Overlapping satellite observations showed a more rapid brightening towards perihelion (Q. Zhang & K. Battams, 2025), but sampled a mixture of gas and dust continuum emission that cannot be meaningfully interpreted; we ignore them here. This activity asymmetry is further corroborated by the fact that the coma of 3I exhibited a significantly steeper surface brightness gradient preperihelion ($\gamma>1$; D. Jewitt et al., 2025; D. Jewitt & J. Luu, 2025) than the shallower postperihelion profile with $\gamma<1$ found in Section III.2 (see Figure 1). The heliocentric index is influenced by many unmeasured parameters of the nucleus, among them the distribution of surface volatiles and the direction of the nucleus spin vector. It is thus impossible to ascribe a unique interpretation to the measured change. We note, however, that measurements of solar system comets show that the heliocentric index can decrease or increase after perihelion with equal probability, and that the steeper post-perihelion index of 3I cannot be regarded as unusual (P. Lacerda et al., 2025).
@@
Finally, we briefly comment on the scattering phase function of 3I. Thanks to the smooth decline in activity and a range of phase angles extending from almost $\sim\!30\arcdeg$ all the way to near-opposition ($\alpha\approx 0\arcdeg$), we were able not only to robustly constrain the linear part of the phase function, but also its opposition surge, neither of which had been previously measured for 3I. The linear portion of the scattering phase function of 3I appears indistinguishable from those of solar system comets. In contrast, to date there remains a persistent dearth of intensity measurements for cometary dust near zero phase angle (V. Rosenbush et al., 2002). Among the few comets that have been studied in this regime, none have exhibited a distinct opposition effect K. J. Meech & D. C. Jewitt (1987). This phenomenon, however, is widely observed among asteroids, planetary rings, and miscellaneous other airless bodies in the solar system (V. Rosenbush et al., 2002). While we acknowledge that the light scattering mechanisms of cometary dust differ from those of airless surfaces, the opposition effect of 3I qualitatively resembles some of the latter. Although K. J. Meech & D. C. Jewitt (1987) found no opposition effect in their selected sample of solar system comets, the brightness enhancement of 3I we observed is broadly in line with their reported upper limit for the opposition surge of $m_{\rm OE}\lesssim 0.2$ mag.
@@
While we cannot rule out the possibility from the available data that the dust grains of 3I are inherently strong backscatterers, we suspect that the detectability of this effect might have been significantly enhanced by concurrent contributions from the orbital plane crossing and the projection of the dust tail. Both factors would increase the total effective cross-section of dust within the photometric aperture, for which we did not corrected in Equation (5). We leave the disentanglement of these competing factors to future dust-modelling work. Taken together with the peculiar polarimetric properties reported by Z. Gray et al. (2025); S. Choi et al. (2026), our results likely suggest that while there are commonalities between dust grains of 3I and solar system comets, unusual physical properties exist.
@@ V Summary
    Using the nucleus extraction technique, we successfully revealed the nucleus of 3I and measured its optical cross-section (the product of geometric albedo and physical cross-section) to be $C_{\rm n}=0.22\pm 0.07$ km2. Adopting a nominal $V$-band geometric albedo of 0.04, typical for cometary nuclei in the solar system, we estimate a nucleus radius of $R_{\rm n}=1.3\pm 0.2$ km. The result is consistent with our independent estimate derived from the reported nongravitational acceleration and production rates of 3I.
@@
    Scattered light by dust grains in the coma displayed a statistically significant opposition surge of $\sim\!0.2$ mag, characterised by an $e$-folding width of $3\arcdeg\pm 1\arcdeg$. This brightening possibly contain coincident contributions from the orbital plane crossing and tail projection. While opposition surges have not been widely reported for solar system cometary dust, the observed brightening is broadly consistent with existing upper limits for solar system comets. The linear portion of the phase function, with a slope of $\beta_{\alpha}=0.026\pm 0.006$ mag degree-1, is indistinguishable from those of solar system comets.
@@
    Based on the discovery statistics of interstellar objects, we estimate a lower limit of $\gtrsim\!1$ 3I-sized (nucleus radius $\sim\!1.3$ km) interstellar object within $r_{\rm H}=4.5$ au at any instant, which is likely a conservative lower bound as inactive interstellar objects of this size would be significantly more difficult to detect. It is likely that comparably bright interstellar objects have passed through the inner solar system in the era of wide-field CCD surveys; we surmise that such objects have been missed.
We thank Alison Vick for her great efforts on scheduling the HST observations to ensure the successful execution of the HST program. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. No NASA funding was used for this work. These observations are associated with program GO/$D$ 18152.

### v2

@@ IV.2 Statistics of Similar Interstellar Interlopers
We attempted to evaluate the statistics of interstellar interlopers similar in size to 3I inside our solar system. Based on 1I/‘Oumuamua alone, D. Jewitt et al. (2017) and A. Do et al. (2018) derived a number density of $\sim\!0.1$-0.2 au-3 for such objects in the solar system. The subsequent discovery of the larger 2I/Borisov was not a fluke but was in line with extrapolated statistics (D. Jewitt et al., 2020). Recently, D. Z. Seligman et al. (2025) reported a number density of $\sim\!3\times 10^{-4}$ au-3 for 3I-like objects with $H_{V}=12.5$. However, because this absolute magnitude led to an overestimate of the nucleus size due to contamination from the coma, the real number density was likely underestimated.
Using Figure 8 from D. Z. Seligman et al. (2025), we estimated the ratio of detectability volume to crossing time for the actual absolute magnitude of 3I’s nucleus ($H_{{\rm n},V}\approx 17.1$; see Section III.2) to be $\sim\!30$ au3 yr-1. The ratio of the ATLAS survey’s detection rate ($\sim\!0.2$ yr-1) to this volume-to-time ratio yields a local number density of $\sim\!7\times 10^{-3}$ au-3 for 3I-sized interstellar objects. Given the discovery of 3I at a heliocentric distance of $r_{\rm H}\approx 4.5$ au, this implies a total number of $\sim\!3$ similar interstellar objects within a sphere of this radius at the Sun at any instant on average.
@@
If extrapolated solely from the statistics based on 1I/‘Oumuamua using a differential size distribution index of -3 for the population of interstellar objects, the expected total number of 3I-sized or larger interstellar objects inside the same volume of space around the Sun will be $\sim\!0.1$-0.3. The likelihood of detecting such an object is given by Poisson statistics to be $\sim\!0.1$-0.2, which drops to $\sim\!\left(1{\rm-}2\right)\%$ for a steeper size distribution of -4. However, these estimates should only be treated as lower limits because cometary activity undoubtedly enhances the detectability of 3I-like objects. We conclude that the discovery of 3I remains in qualitative agreement with the previous statistics of the interstellar object population.
The arrival rate of 3I-like interstellar objects, estimated from the ratio of their total number within the 4.5 au radius sphere to the corresponding crossing timescale, $\sim\!2\times 10^{7}$ s, is $\sim\!4$ yr-1 assuming a steady-state population. We found that the maximum crossing time for this distance would be $\sim\!1\times 10^{8}$ s, corresponding to objects on barely hyperbolic trajectories, which implies a lower limit on the arrival rate of $\gtrsim\!0.6$ yr-1. A qualitatively similar result for the detection rate of 3I-sized objects, $\sim\!0.3$-0.6 yr-1, can be reached by scaling from the relation presented in D. J. Hoover et al. (2022) using a differential size distribution index of -3, whereas a steeper index of -4 would yield a rate in the range of $\sim\!0.02$-0.04 yr-1, smaller by an order of magnitude.
@@
Although all-sky surveys such as Pan-STARRS and Catalina have been operational for over a decade, only two other interstellar objects were previously discovered, both of which were visually fainter than 3I. Around its peak brightness, 3I was sufficiently bright that even visual comet hunters could have detected it had it passed through the inner solar system prior to the era of automated sky surveys. Here we simplistically estimate the likelihood that no 3I-like interstellar objects passed by between mid-1990s, the advent of modern sky surveys, and the discovery of 1I/‘Oumuamua in 2017. Over this $\sim\!20$-year interval, the expected number of detections is $\sim\!4$, assuming a detection rate of $\sim\!0.2$ yr-1. According to Poisson statistics, the probability of no detectable objects in this period is $\sim\!\exp\left(-4\right)\approx 2\%$. Repeating this order-of-magnitude estimation by scaling from the relation in D. J. Hoover et al. (2022) with an index of -3 yields a probability of $\lesssim\!3\times 10^{-3}$, which is negligible. However, the probability of a null detection increases considerably to 0.5-0.6 if a steeper index of -4 is adopted. While earlier surveys were undeniably less sensitive if compared to contemporary ones, they were unlikely to be less efficient by many orders of magnitude. Therefore, if the size distribution of the interstellar object population is not steep, it appears highly probable that several 3I-like interstellar objects passed through the inner solar system undetected between the advent of automated sky surveys and the discovery of 1I/‘Oumuamua. Alternatively, the lack of prior detections may imply a steep size distribution for the population, potentially in agreement with the inference in D. Jewitt et al. (2017).
@@ IV.3 Dust Activity & Scattering Phase Function
The use of an alternative phase function model yielded a mean activity index of $\bar{n}=4.5\pm 0.03$, which is not statistically distinguishable from the one with the linear model, implying that the result of the activity index is robust and relatively independent of the chosen phase correction. While preperihelion continuum photometry from $r_{\rm H}=4.6$ au to 1.8 au revealed a steady brightening of 3I well described by an activity index $n=3.8\pm 0.3$ (D. Jewitt & J. Luu, 2025), our postperihelion result implies that, over similar ranges of heliocentric distance, 3I faded more rapidly on the outbound leg than it brightened on the inbound leg. [^4] Overlapping satellite observations showed a more rapid brightening towards perihelion (Q. Zhang & K. Battams, 2025), but sampled a mixture of gas and dust continuum emission that cannot be meaningfully interpreted; we ignore them here. This activity asymmetry is further corroborated by the fact that the coma of 3I exhibited a significantly steeper surface brightness gradient preperihelion ($\gamma>1$; D. Jewitt et al., 2025; D. Jewitt & J. Luu, 2025) than the shallower postperihelion profile with $\gamma<1$ found in Section III.2 (see Figure 1). The heliocentric index is influenced by many unmeasured parameters of the nucleus, among them the distribution of surface volatiles and the direction of the nucleus spin vector. It is thus impossible to ascribe a unique interpretation to the measured change. We note, however, that measurements of solar system comets show that the heliocentric index can decrease or increase after perihelion with equal probability, and that the steeper post-perihelion index of 3I cannot be regarded as unusual (P. Lacerda et al., 2025).
@@
Finally, we briefly comment on the scattering phase function of 3I. Thanks to the smooth decline in activity and a range of phase angles extending from almost $\sim\!30\arcdeg$ all the way to near-opposition ($\alpha\approx 0\arcdeg$), we were able not only to robustly constrain the linear part of the phase function, with a slope of $\bar{\beta}_{\alpha}=0.026\pm 0.006$ mag deg-1, but also its opposition surge, neither of which had been previously measured for 3I. The linear portion of the scattering phase function of 3I appears indistinguishable from those of solar system comets. In contrast, to date there remains a persistent dearth of intensity measurements for cometary dust near zero phase angle (V. Rosenbush et al., 2002). Among the few comets that have been studied in this regime, none have exhibited a distinct opposition effect K. J. Meech & D. C. Jewitt (1987). This phenomenon, however, is widely observed among asteroids, planetary rings, and miscellaneous other airless bodies in the solar system (V. Rosenbush et al., 2002). While we acknowledge that the light scattering mechanisms of cometary dust differ from those of airless surfaces, the opposition effect of 3I qualitatively resembles some of the latter. Although K. J. Meech & D. C. Jewitt (1987) found no opposition effect in their selected sample of solar system comets, the brightness enhancement of 3I we observed is broadly in line with their reported upper limit for the opposition surge of $m_{\rm OE}\lesssim 0.2$ mag.
@@
While we cannot rule out the possibility from the available data that the dust grains of 3I are inherently strong backscatterers, we suspect that the detectability of this effect might have been significantly enhanced by concurrent contributions from the orbital plane crossing and the projection of the dust tail. Both factors would increase the total effective cross-section of dust within the photometric aperture, for which we did not correct in Equation (5). We leave the disentanglement of these competing factors to future dust-modelling work. Taken together with the peculiar polarimetric properties reported by Z. Gray et al. (2025); S. Choi et al. (2026), our results suggest that while there are commonalities between dust grains of 3I and solar system comets, unusual physical properties exist.
@@ V Summary
    Using the nucleus extraction technique, we successfully revealed the nucleus of 3I and measured its optical cross-section (the product of geometric albedo and physical cross-section) to be $C_{\rm n}=0.22\pm 0.07$ km2. Adopting a nominal $V$-band geometric albedo of 0.04, typical for cometary nuclei in the solar system, we estimate an effective nucleus radius of $R_{\rm n}=1.3\pm 0.2$ km. The result is consistent with our independent estimate derived from the reported nongravitational acceleration and production rates of 3I.
@@
    Scattered light by dust grains in the coma displayed a statistically significant opposition surge of $\sim\!0.2$ mag, characterised by an $e$-folding width of $3\arcdeg\pm 1\arcdeg$. This brightening possibly contain coincident contributions from the orbital plane crossing and tail projection. While opposition surges have not been widely reported for solar system cometary dust, the observed brightening is broadly consistent with existing upper limits for solar system comets. The linear portion of the phase function, with a slope of $\bar{\beta}_{\alpha}=0.026\pm 0.006$ mag degree-1, is indistinguishable from those of solar system comets.
@@
    Based on the discovery statistics of interstellar objects, we estimate a minimum of three 3I-sized (nucleus radius $\sim\!1.3$ km) interstellar object within $r_{\rm H}=4.5$ au at any given instant, which is likely a conservative lower bound as inactive interstellar objects of this size would be significantly more difficult to detect than an active object like 3I. It is probable that comparably bright interstellar objects have passed through the inner solar system during the era of wide-field CCD surveys; we surmise that such objects have been missed. Alternatively, the lack of prior detections may imply a steep size distribution for the interstellar object population.
We thank Alison Vick for her great efforts in scheduling the HST observations and ensuring the successful execution of the HST program. We also thank Darryl Seligman and Maria Womack for their prompt and insightful reports, which was helpful in improving our manuscript. This research is based on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5–26555. No NASA funding was used for this work. These observations are associated with program GO/$D$ 18152.